Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:

Direct materials: 5 pounds at $10 per pound $ 50
Direct labor: 3 hours at $17 per hour 51
Variable overhead: 3 hours at $7 per hour 21
Total standard cost per unit $ 122

The planning budget for March was based on producing and selling 24,000 units. However, during March the company actually produced and sold 30,600 units and incurred the following costs:

  1. Purchased 170,000 pounds of raw materials at a cost of $9.00 per pound. All of this material was used in production.
  2. Direct laborers worked 68,000 hours at a rate of $18 per hour.

  3. Total variable manufacturing overhead for the month was $512,040.

Questions:

6. If Preble had purchased 183,000 pounds of materials at $9.00 per pound and used 170,000 pounds in production, what would be the materials quantity variance for March?

7. What direct labor cost would be included in the company’s planning budget for March?

8. What direct labor cost would be included in the company’s flexible budget for March?

9. What is the labor rate variance for March?

10. What is the labor efficiency variance for March?
Answer #1

6) Material quantity variance = (30600*5-170000)*10 = 170000 U

7) Labor in planning budget = 24000*51 = 1224000

8) Labor in flexible budget = 30600*51 = 1560600

9) Labor rate variance = (17-18)*68000 = 68000 U

10) Labor efficiency variance = (30600*3-68000)*17 = 404600 F